OpenStudy (anonymous):

What is the volume of a cube with a side length of 4 inches?

OpenStudy (anonymous):

4 cubed = 64

OpenStudy (solomonzelman):

all sides of cube are equal and the formula for volume of cube (where 's" is side) \(\huge\color{blue}{ \sf V=s^3 }\) i your case. \(\huge\color{blue}{ \sf V=4^3=64 }\)